7/9 -(-13/18)= -1/18
Answer by stanbon(75887)

7/9 -(-13/18)= -1/18
I don't know where you got this equation but
it is not true.
-----------
14/18 +13/18 = 27/18
==========================
Cheers,
Stan H.